# Renormalization of Interacting Random Graph Models

### v2

@@ I.2 Universality classes of graph Hamiltonians
Truncating the effective Hamiltonian to bilinear interactions might seem like an academic oversimplification at first glance. However, as elaborated in the next section, Eq. 2 can also be viewed as the Hamiltonian of an equivalent disordered spin system with couplings of arbitrary order (i.e. in terms of degrees of nearest neighbors), whose effective dimensionality is determined by the maximum coordination number encoded in the $h^{(2)}_{ijkl}$, and presuming a hierarchy for the latter over higher order couplings. Consequently, the result of successive RG transformations for a large class of bare effective Hamiltonians would be to flow towards the form Eq. 9, which therefore represents a universality class of equivalent random graph models, and therefore an important object of study on its own terms. Moreover, from a practical and application intensive point of view, Eq. 9 can also effectively model temporal Markov network processes, where RG transformations would correspond to temporal coarse graining. We elaborate on this in more detail in an accompanying study [36], and commence our formal study of random graph renormalization in the following sections. We conclude our investigation with a discussion of the implications of our findings, the most pertinent of which is that within the universality class of bilinear graph Hamiltonians, all models are infra-red trivial, meaning that all pairwise condition flow to irrelevance at large enough scales. We translate the ramifications of this for a number of applications, and close with a discussion of open questions and future directions.
@@ II Line graphs, spin chains, and closed RG transformations
where we’ve relabeled the indices to run over the integers again after each successive uniform decimation.
@@
Figure 1: Schematic rendering of maximum coordination two couplings (top) and a single bare coordination number three coupling (bottom). More generally, successive RG steps increase the maximum coordination number, and flow towards all-to-all couplings (bottom most diagram).
@@ III Homogeneous RG flow
A simple but informative case to get our bearings with is the setting where each of the $\{g_{a},\varepsilon_{a}\}$ couplings in Eq. 12 are respectively fixed to be the same, and we successively decimate, or marginalize over alternate links on the line graph. In this case, Eqs. 14 reduce to the following system of equations:
@@
Figure 2: RG flow for homogeneous coordination number two interactions towards an Erdős-Renyi fixed line. Left panel: A shearing behavior for positively correlated conditioning is apparent, along with a separatrix along the line $\varepsilon=-g$ for $g>0$. This corresponds to the vanishing of the beta function of external magnetic field $h:=\varepsilon+g$ along the line $h\equiv 0$ for the analog spin glass system, illustrated by the right panel.
The various features of the flow diagram lend themselves to straightforward interpretation. We first note that the part of the diagram below the $g=0$ axis is only meaningful for a single RG step. The reason is that implicit in Eq. 16 is the fact that were we to begin with a negative value for the bare $g$, a single decimation iteration will return a positive value for $g$. Although the structure is far richer when we turn our attention to inhomogeneous and disordered flows, the latter feature of considering only homogeneous interactions can be readily understood from the fact that on the analog spin system, negative values of $g$ corresponds to uniform anti-ferromagnetic couplings. When one considers the effect of marginalizing over two adjacent anti-ferromagnetic couplings, the result is a effective ferromagnetic coupling (i.e. $\widetilde{g}>0$) that maintains its ferromagnetic nature under further RG iterations.
@@ IV.1 Sparse disorder
Inserting these into Eqs. IV and IV yield the following induced renormalization group transformations on the probability assignments upon iteration:
@@ V.1 Peer, preferential, susceptibility, and reinforcement effects on random networks
Such terms would correspond to terms in an effective Hamiltonian of the general form Eq. 10 as opposed to Eq. 12 on the corresponding line graph representation. Non-liner attachment kernels of the form $p_{i}\propto f(k_{i})$ would necessitate the addition of a series of higher order interactions of the form
@@ Appendix A Generating functionals, and when renormalization $\equiv$ marginalization
Consider now the situation where we have a partition of our set of random variables into two disjoint subsets $X=\{\textbf{y},\textbf{x}\}$, where without loss of generality, we order them such that $\{\textbf{y}\}:=\{x_{1},...,x_{k}\}$ and $\{\textbf{x}\}:=\{x_{k+1},...,x_{\mathcal{N}}\}$, with joint probability distribution given by $\Pi(\textbf{y},\textbf{x})$. We furthermore presume that we wither have no means of accessing or measuring the x degrees of freedom, or we are simply uninterested in computing any statistical moments that involve them. In that case, any observable $\mathcal{O}(\textbf{y})$ that is constructed exclusively out of the y variables can be re-expressed as
@@ Appendix C Induced RG flow as time reversed drift-diffusion
Although many interesting structures relating to parameter relevance and renormalization group flow can be inferred from and for the Fisher-Rao metric alone [^17] In addition to the direct correspondence between the eigenvalues of the Fisher-Rao metric and RG parameter relevance discussed earlier [24], one can also show that the metric undergoes a geometric flow under coarse graining induced by the beta function vector field [79]. The information metric also geometrizes the study of the phase structure and critical exponents of a statistical mechanical system, as can be seen via Eqs. 74 and C, so that $G_{ab}(\theta)=-\partial_{a}\partial_{b}\log Z$. Recognizing the latter as the generalized free energy, we see immediately that singularities on the statistical manifold identify phase transitions [80, 81, 82], and allows one to define and compute distances between theories [83]., it will plays only a spectator role in what follows.
@@
At this stage, it behooves us to be more precise about the disorder probability measure. The freedom to reparameterize the disorder variables (as we have done freely in the previous sections), which is to allow for arbitrary coordinate transformations on $\mathcal{M}$, would require us to work covriantly and incorporate a volume factor $\sqrt{G}$ (with $G$ being the determinant of $G_{ab}$) into the measure, with all derivatives corresponding to covariant derivatives with metric compatible connections. Although the derivation that follows proceeds analogously for arbitrary connections, we restrict ourselves to flat connections over a specific coordinatization such that the measure is as indicated in Eqs. 80 and 82. This is not just to simplify the discussion, but also because a more comprehensive treatment would also admit more general possibilities for connections on the statistical manifold (e.g. the so-called one parameter family of $\alpha$-connections [27, 28]), which would deserve a separate discussion [^20] In the present context, the probability assignments Eq. 74 correspond to an exponential family of distributions in the information geometric framework, and our choice to use the flat metric connections corresponds to working with an $\alpha$-connection with $\alpha=1$ [28, 27].

### v3

@@ I.2 Universality classes of graph Hamiltonians
Truncating the effective Hamiltonian to bilinear interactions might seem like an academic oversimplification at first glance. However, as elaborated in the next section, Eq. 2 can also be viewed as the Hamiltonian of an equivalent disordered spin system with couplings of arbitrary order (i.e. in terms of degrees of nearest neighbors), whose effective dimensionality is determined by the number of nearest neighbors each site has. We refer to this quantity as the coordination number of a link, which is encoded in all the $h^{(2)}_{ijkl}$ terms being different from zero. Were one to presume a hierarchy for these terms over higher order couplings, the result of successive RG transformations for a large class of bare effective Hamiltonians would be to flow towards the form Eq. 9, which therefore represents a universality class of equivalent random graph models, and therefore an important object of study on its own terms. Moreover, from a practical and application intensive point of view, Eq. 9 can also effectively model temporal Markov network processes, where RG transformations would correspond to temporal coarse graining. We elaborate on this in more detail in an accompanying study [36], and commence our formal study of random graph renormalization in the following sections.
I.3 Outline
The goal of this paper is to set up a general framework through which one might quantify the scaling of interactions in random graph models. That is, to determine the relevance, or irrelevance of the conditionings of different graph variables on each other. Although this is highly non-trivial problem in full generality, we focus in what follows on a restricted class of transformations for which closed form renormalization group transformations are available, which we argue to represent a universality class of graph Hamiltonians.
In the sections that follow, we first set up the formalism of random graph renormalization in the line graph representation in section II, elaborating on the existence of closed form RG transformations for maximum coordination number two pairwise interactions. We then consider how these interactions renormalize when the couplings are homogeneously assigned in III, and generalize to the disordered setting in IV, where we immediately find that the problem can be recast in terms of determining the the induced flow on the disorder probability assignments. We understand this flow to be akin to time reversed drift diffusion, as detailed in appendix C, with additional details relevant to our discussion collected in appendices A B. We conclude our investigation with a discussion of the implications of our findings, the most pertinent of which is that within the universality class of bilinear graph Hamiltonians, all models are infra-red trivial, meaning that all pairwise conditions flow to irrelevance at large enough scales. We translate the ramifications of this for a number of applications, and close with a discussion of open questions and future directions.
@@ II Line graphs, spin chains, and closed RG transformations
where we’ve relabeled the indices to run over the integers again after each successive uniform decimation. A simple derivation of these equations can be found in the context of the Ising Model with sparse disorder (see e.g. [38]), or in our accompanying paper [36].
@@
Figure 1: Schematic rendering of maximum coordination two couplings (top) and a single bare coordination number three coupling (bottom) under RG transformations. The red crosses indicate the sites that are decimated at each RG step. More generally, successive RG steps increase the maximum coordination number, and flow towards all-to-all couplings (bottom most diagram).
@@ III Homogeneous RG flow
A simple but informative case to get our bearings with is the setting where each of the $\{g_{a},\varepsilon_{a}\}$ couplings in Eq. 12 are respectively fixed to be the same, and we successively decimate, or marginalize over alternate links on the line graph (see for a thorough derivation [36]). In this case, Eqs. 14 reduce to the following system of equations:
@@
Figure 2: RG flow for homogeneous coordination number two interactions towards an Erdős-Renyi fixed line, where arrows indicate direction of flow and color determine the its intensity, lighter being stronger. Upper panel: A shearing behavior for positively correlated conditioning is apparent, along with a separatrix along the line $\varepsilon=-g$ for $g>0$. This corresponds to the vanishing of the beta function of external magnetic field $h:=\varepsilon+g$ along the line $h\equiv 0$ for the analog spin glass system, illustrated by the bottom panel.
The various features of the flow diagram lend themselves to straightforward interpretation. We first note that the part of the diagram below the $g=0$ axis is only meaningful for a single RG step. The reason is that implicit in Eq. 16 is the fact that were we to begin with a negative value for the bare $g$, a single decimation iteration will return a positive value for $g$. Although the structure is far richer when we turn our attention to inhomogeneous and disordered interactions, the latter feature of considering only homogeneous interactions can be readily understood from the fact that on the analog spin system, negative values of $g$ corresponds to uniform anti-ferromagnetic couplings. When one considers the effect of marginalizing over two adjacent anti-ferromagnetic couplings, the result is a effective ferromagnetic coupling (i.e. $\widetilde{g}>0$) that maintains its ferromagnetic nature under further RG iterations.
@@ IV.1 Sparse disorder
where $\Omega(x_{n})$ and $\Pi(y_{n})$ are the bare distributions of the $x$ and $y$ variables for the site where the disorder is localized. Inserting these into Eqs. IV and IV yield the following induced renormalization group transformations on the probability assignments upon iteration:
@@ V.1 Peer, preferential, susceptibility, and reinforcement effects on random networks
Such terms would correspond to terms in an effective Hamiltonian of the general form Eq. 10 as opposed to Eq. 12 on the corresponding line graph representation. Non-linear attachment kernels of the form $p_{i}\propto f(k_{i})$ would necessitate the addition of a series of higher order interactions of the form
@@ Appendix A Generating functionals, and when renormalization $\equiv$ marginalization
Consider now the situation where we have a partition of our set of random variables into two disjoint subsets $X=\{\textbf{y},\textbf{x}\}$, where without loss of generality, we order them such that $\{\textbf{y}\}:=\{x_{1},...,x_{k}\}$ and $\{\textbf{x}\}:=\{x_{k+1},...,x_{\mathcal{N}}\}$, with joint probability distribution given by $\Pi(\textbf{y},\textbf{x})$. We furthermore presume that we either have no means of accessing or measuring the x degrees of freedom, or we are simply uninterested in computing any statistical moments that involve them. In that case, any observable $\mathcal{O}(\textbf{y})$ that is constructed exclusively out of the y variables can be re-expressed as
@@ Appendix C Induced RG flow as time reversed drift-diffusion
Although many interesting structures relating to parameter relevance and renormalization group flow can be inferred from and for the Fisher-Rao metric alone [^17] In addition to the direct correspondence between the eigenvalues of the Fisher-Rao metric and RG parameter relevance discussed earlier [24], one can also show that the metric undergoes a geometric flow under coarse graining induced by the beta function vector field [79]. The information metric also geometrizes the study of the phase structure and critical exponents of a statistical mechanical system, as can be seen via Eqs. 74 and C, so that $G_{ab}(\theta)=-\partial_{a}\partial_{b}\log Z$. Recognizing the latter as the generalized free energy, we see immediately that singularities on the statistical manifold identify phase transitions [80, 81, 82], and allows one to define and compute distances between theories [83]., it will play only a spectator role in what follows.
@@
At this stage, it behooves us to be more precise about the disorder probability measure. The freedom to reparameterize the disorder variables (as we have done freely in the previous sections), which is to allow for arbitrary coordinate transformations on $\mathcal{M}$, would require us to work covariantly and incorporate a volume factor $\sqrt{G}$ (with $G$ being the determinant of $G_{ab}$) into the measure, with all derivatives corresponding to covariant derivatives with metric compatible connections. Although the derivation that follows proceeds analogously for arbitrary connections, we restrict ourselves to flat connections over a specific coordinatization such that the measure is as indicated in Eqs. 80 and 82. This is not just to simplify the discussion, but also because a more comprehensive treatment would also admit more general possibilities for connections on the statistical manifold (e.g. the so-called one parameter family of $\alpha$-connections [27, 28]), which would deserve a separate discussion [^20] In the present context, the probability assignments Eq. 74 correspond to an exponential family of distributions in the information geometric framework, and our choice to use the flat metric connections corresponds to working with an $\alpha$-connection with $\alpha=1$ [28, 27].